REGRESSIONAL CHARACTERIZATION OF THE GENERALIZED INVERSE GAUSSIAN POPULATION

J. PUSZ

Mathematical Institute, Warsaw University of Technology, Plac Politechniki 1, 00-661 Warsaw, Poland

(Received January 22, 1996; revised July 19, 1996)

Abstract. A characterization of the generalized inverse Gaussian population is obtained from a condition of constant regression of a suitable statistic on a linear one.

Key words and phrases: Generalized inverse Gaussian distribution, regression, characterization.